OpenStudy (anonymous):

value of cos^2x(1+cos^2x) where sinx=cos^2x

OpenStudy (kc_kennylau):

When you are asked to find a value of a particular thing without any number given, you usually would try to form patterns involving \(\cos^2\theta+\sin^2\theta\) because its value is 1.

OpenStudy (anonymous):

The information you are given put that into the equation and see if u can mix the equation up to your own good and finding the answer Hint:- Recall the Fundamental identities

OpenStudy (kc_kennylau):

If I were you, I would expand it first. (expand = distribute)

OpenStudy (anonymous):

YEah expand it!

OpenStudy (anonymous):

yup